The width of a retangle is 15 feet less than its length, and the perimeter is 210 feet. Can you show me how to find the width and the length.
Answer by venugopalramana(3286)

The width ( W SAY) of a retangle is 15 feet less than its length( L SAY), and
W =L-15

the perimeter( P SAY) is 210 feet.
P= 2L + 2W
210 = 2L +2(L-15)=2L+2L-30=4L-30
4L-30=210
4L=210+30=240
L=240/4 = 60
W = 60-15 =45


Can you show me how to find the width and the length.